OpenStudy (anonymous):

the distance of a particle from a fixed point O is given by s= 3cos2t + 4sin2t show that the velocity v and the acceleration a are given by v^2 + 4s^2 =100, a+4s=0. Please give me a hint someone!!

OpenStudy (anonymous):

Velocity is the derivative of the position and acceleration is the second derivative of position and first derivative of velocity!

OpenStudy (anonymous):

also @stamp has a weak pellet

OpenStudy (anonymous):

I know that but what does the v^2 + 4s^2 =100 mean?

OpenStudy (anonymous):

cause when I diffrentiate I get trig ratios.

OpenStudy (anonymous):

I don't have v , I only have s !!

OpenStudy (anonymous):

no read the question ,

OpenStudy (anonymous):

it says

OpenStudy (anonymous):

he distance of a particle from a fixed point O is given by s= 3cos2t + 4sin2t show that the velocity v and the acceleration a are given by v^2 + 4s^2 =100, a+4s=0.

OpenStudy (anonymous):

Task 1, find the velocity equation. Can you do it?

OpenStudy (anonymous):

Can you take the derivative?

OpenStudy (anonymous):

yes is it -6sint+8cos2t?

OpenStudy (anonymous):

Okay, so then substitute that in for \(v\) in \(v^2+4s^2=100\)

OpenStudy (anonymous):

\[ v^2+4s^2=(-6\sin t+8\cos2t)^2+4(3\cos2t + 4\sin2t)^2=\dots \]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Hint: \(\sin^2(t)+\cos^2(t)=1\)

OpenStudy (anonymous):

Also actually you got velocity wrong I think, should have a \(2t\) rather than just \(t\).

OpenStudy (anonymous):

oh yes ! sorry